OpenStudy (anonymous):

find the area between the curve y=abs. vale of lnx and the x-axis, and lines x=e and x=1/e

OpenStudy (anonymous):

its that hard?

OpenStudy (anonymous):

\[\int_{\frac{1}{e}}^1 -\log (x) \, dx+\int_1^e \log (x) \, dx\]

OpenStudy (anonymous):

why negative log?

OpenStudy (anonymous):

the change?

OpenStudy (anonymous):

because log[x] is negative between 1/e and 1, we want positive

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so what is the integral of - log x?

OpenStudy (anonymous):

i dont know how to do it thou

OpenStudy (anonymous):

Try graphing you will see

OpenStudy (anonymous):

i did.....

OpenStudy (anonymous):

http://tinyurl.com/3kmg2xq

OpenStudy (anonymous):

This is how log(x) looks like http://tinyurl.com/3ecc87y

OpenStudy (anonymous):

but how do i do the integral?